# On continuity of state-dependent utilities

### v1

@@ Abstract
State-dependent preferences for a general Savage’s state space were shown in Wakker and Zank (1999) to admit a numerical representation in the form of the integral of a state-dependent utility, as soon as pointwise continuity of the preference ordering is assumed. In this paper we prove that such a state-dependent function inherits pointwise continuity from the preference ordering, providing in this way a positive answer to a conjecture posed in the aforementioned seminal work.
We further apply this result to obtain an explicit representation of conditional Chisini means in the form of a conditional certainty equivalent.
keywords:
@@
pointwise continuity, state-dependent expected utility, conditional certainty equivalent, nonlinear expectation
MSC:
[2020] 91B06, 91B08, 60A05
@@
[…]
1 Introduction
For a general state space $(\Omega,{\mathcal{F}})$, Wakker and Zank WZ99 provide a characterization of those preference relations $\succeq$ on the space of bounded random acts which admit a numerical representation through an additively decomposable functional (WZ99 , Theorem 11). This result extends to an infinite dimensional setup the theory developed by Debreu De60 . By adopting the stronger notion of pointwise continuity of $\succeq$, Theorem 11 in WZ99 can be specialized to the following integral representation form.
Theorem 1.1 (WZ99 , Theorem 12 and CL06 , Theorem 5)
Assume that ${\mathcal{F}}$ contains at least three disjoint non-null events. Then the following two statements are equivalent:
  1. (i)
    There exists a countably additive finite measure $\mu$ on $\Omega$ and a function $\mathbb{u}(\omega,\cdot):\mathbb{R}\to\mathbb{R}$ strictly increasing for every $\omega\in\Omega$, such that $\succeq$ is represented by the pointwise continuous integral
    $f\to\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mu(\omega).$
  2. (ii)
    $\succeq$ is a preference order satisfying strict monotonicity, the sure-thing principle and pointwise continuity  [^1] See page (SM) below for further details on the preference axioms involved.
The previous theorem leaves an interesting issue unsolved as declared by the authors WZ99 , page 29:
“[…] Pointwise continuity is not overly restrictive because continuity of each state-dependent utility implies pointwise continuity. Whether the reversed implication holds, i.e., whether continuity of the functions $\mathbb{u}$, in Statement (i) of Theorem 12 holds (after appropriate modification on null events), is an open question to us. ”
To the best of our knowledge this question did not find an answer in the subsequent literature. Among the principal aims of the present paper, Theorem 3.1 stated below provides the construction of a pointwise continuous version of the state dependent utility $\mathbb{u}$. Such a result has the immediate benefit of clarifying the role of continuity in the application of the Theorem 1.1. For example, in Appendix A of MM21 a somewhat cumbersome notion of continuity is introduced to overcome some technical hurdles in the proof of the main result. Guaranteeing the existence of a continuous $\mathbb{u}$ allows to circumvent issues of this kind.
As an important byproduct, we provide a representation of the conditional Chisini mean $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ previously introduced and studied in DM23 . The conditional Chisini mean is a nonlinear extension of the standard concept of conditional expectation, which takes inspiration from the original Chisini’s intuition Ch29 that the notion of “expected value” should be introduced as the solution of a functional equation. For a sub-$\sigma$-algebra ${\mathcal{G}}$ of ${\mathcal{F}}$, representing intermediate information available, we show that $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)=\mathbb{u}_{\mathcal{G}}^{-1}(\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right])$, where $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$ is the conditional expectation of the random outcome $\omega\mapsto\mathbb{u}(\omega,f(\omega))$ and $\mathbb{u}_{\mathcal{G}}(\omega,x)=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(x)\middle|{\mathcal{G}}\right](\omega)$ is the ${\mathcal{G}}$-measurable projection of the state dependent utility $\mathbb{u}$. This representation can be easily compared to the general formulation proved in CVMM , Lemma 5.2, under additional law invariance assumptions, which traces back to the celebrated Nagumo, de Finetti, Kolmogorov Theorem (see HLP52 for a review). It is also worth mentioning that in this case, $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ takes the form of a conditional certainty equivalent based on state-dependent utilities (see FM11cce ).
2 Notations and preliminaries
[…]
Whenever working with the product space $\Omega\times\mathbb{R}$, we shall make use of the product $\sigma$-algebra ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}$ i.e. the $\sigma$-algebra generated by elements of the form $F\times B$ with $F\in{\mathcal{F}}$ and $B\in\mathbb{B}_{\mathbb{R}}$. Given a probability measure $\mathbb{P}$, $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space. A (finite) signed measure $\nu$ on $(\Omega,{\mathcal{F}})$ is said to be dominated by $\mathbb{P}$ ($\nu\ll\mathbb{P}$) if, for any $A\in{\mathcal{F}}$, $\mathbb{P}(A)=0$ implies $\lvert\nu\rvert(A)=0$, where $\lvert\nu\rvert$ is the total variation measure of $\nu$. By the Hahn decomposition theorem, $\nu\ll\mathbb{P}$ is equivalent to: $\mathbb{P}(A)=0$ implies ${\nu}(A)=0$.
For a given probability $\mathbb{P}$ on $(\Omega,{\mathcal{F}})$ we shall denote the space of integrable functions by ${\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})=\left\{f\in{\mathcal{L}}(\Omega,{\mathcal{F}}):\mathbb{E}_{\mathbb{P}}\left[|f|\right]<\infty\right\}$.
For any $f\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ and a sub $\sigma$-algebra ${\mathcal{G}}$, $\mathbb{E}_{\mathbb{P}}\left[f\mid{\mathcal{G}}\right]$ represents the conditional expectation i.e.
@@ 2 Notations and preliminaries
In the following we shall assume familiarity with the properties of conditional expectation. We refer the reader to Williams , page 88, for a comprehensive review of such properties. A preference relation is a binary relation $\succeq$ on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ which satisfies completeness and transitivity and induces a (strict) preference order $\succ$ by $f\succ g$ if and only if $f\succeq g$ but $g\nsucceq f$. In addition we set $g\sim f$ whenever $g\succeq f$ and $f\succeq g$. We denote
@@
$\succeq$ satisfies the sure-thing principle whenever, considering arbitrary $f,g,h\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and $A\in{\mathcal{F}}$ such that $f\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}\succeq g\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}$, for every $\tilde{h}\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we have $f\mathbf{1}_{A}+\tilde{h}\mathbf{1}_{A^{c}}\succeq g\mathbf{1}_{A}+\tilde{h}\mathbf{1}_{A^{c}}$;
@@
Assumptions (SM), (ST) and (PC) represent the minimal framework we will be working with throughout the paper. We recall that, provided ${\mathcal{F}}$ contains at least three distinct non-null events, by Theorem 1.1 any preference ordering satisfying these assumptions admits an integral representation.
3 Statement of the main results
We are now ready to state the two main results of this paper. The first one precisely answers to the question posed by WZ99 as announced in the introduction and from this perspective it provides an important addendum to Theorem 1.1.
Theorem 3.1

### v2

# On conditioning and consistency for nonlinear functionals

@@ Abstract
We consider a family of conditional nonlinear expectations defined on the space of bounded random variables and indexed by the class of all the sub-sigma-algebras of a given underlying sigma-algebra. We show that if this family satisfies a natural consistency property, then it collapses to a conditional certainty equivalent defined in terms of a state-dependent utility function. This result is obtained by embedding our problem in a decision theoretical framework and providing a new characterization of the Sure-Thing Principle. In particular we prove that this principle characterizes those preference relations which admit consistent backward conditional projections. We build our analysis on state-dependent preferences for a general state space as in Wakker and Zank (1999) and show that their numerical representation admits a continuous version of the state-dependent utility. In this way, we also answer positively to a conjecture posed in the aforementioned paper.
@@
Keywords: Sure-Thing Principle, pointwise continuity, state-dependent expected utility, conditional certainty equivalent, nonlinear expectation, time consistency, Chisini mean.
1 Introduction
In the last decades an intensive interplay between Decision Theory and Financial Mathematics led to the development of stochastic portfolio optimization in financial markets, started in the seminal contribution of Merton [18] and later followed by a flourishing stream of literature. The maximization of an economic agent’s utility can be formulated as a stochastic control problem adopting a preference ordering $\succeq$ on the space of random variables. Such a preference is usually represented as an expected utility and therefore automatically satisfies the so-called Sure-Thing Principle.
The Sure-Thing Principle (commonly referred to as Axiom P2) began to enjoy popularity after the celebrated book by L.J. Savage [21], where it was illustrated by the following (to some extent tautological) example
A businessman contemplates buying a certain piece of property. He considers the outcome of the next presidential election relevant. So, to clarify the matter to himself, he asks whether he would buy if he knew that the Democratic candidate were going to win, and decides that he would. Similarly, he considers whether he would buy if he knew that the Republican candidate were going to win, and again finds that he would. Seeing that he would buy in either event, he decides that he should buy, even though he does not know which event obtains, or will obtain, as we would ordinarily say. [21, page 21].
We can rephrase the example as follows: if a decision maker prefers a position $f$ to $g$, either knowing that the event $B$ or $B^{C}$ (the complement) has occurred, then she should prefer $f$ to $g$ even if she knows nothing about $B$. As depicted in [13, page 98] “P2 is akin to requiring that you have conditional preferences, namely, preferences between $f$ and $g$ conditional on $A$ occurring, and that these conditional preferences determine your choice between $f$ and $g$ if they are equal in case $A$ does not occur”. In this paper we explore how this principle is a strongly characterizing property for the existence of conditional (time consistent) backward projections of the initial ordering relation. The validity of the Sure-Thing Principle has been questioned and highly debated as a consequence of the Ellsberg Paradox, (see e.g. [13, Section 12.3]). From this perspective our findings might be seen as arguing the somewhat excessive strength of this axiom in dynamic decision making.
We work with a preference $\succeq$ on the space of bounded random variables $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ which satisfies monotonicity and pointwise continuity. We show that for this class of preferences the Sure-Thing Principle is equivalent to the existence of a ${\mathcal{G}}$-measurable $\succeq$-equivalent (which we shall refer to as conditional Chisini mean [11]) for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and ${\mathcal{G}}$ being a $\sigma$-algebra contained in ${\mathcal{F}}$ (see Theorem 4 below). The conditional Chisini mean is a nonlinear extension of the standard concept of conditional expectation, which takes inspiration from the original Chisini’s intuition [7] that the notion of “expected value” should be introduced as the solution of a functional equation.
Additionally we provide in Theorem 8 a representation of the conditional Chisini mean in the form $\mathbb{u}_{\mathcal{G}}^{-1}(\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right])$, where $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$ is the conditional expectation of the random outcome $\omega\mapsto\mathbb{u}(\omega,f(\omega))$ and $\mathbb{u}_{\mathcal{G}}(\omega,x)=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(x)\middle|{\mathcal{G}}\right](\omega)$ is the ${\mathcal{G}}$-measurable projection of the state-dependent utility $\mathbb{u}$. This representation can be easily compared to the general formulation proved in [6], Lemma 5.2, under additional law invariance assumptions, which traces back to the celebrated Nagumo, de Finetti, Kolmogorov Theorem (see [14] for a review). It is also worth mentioning that Theorem 8 leads to a theoretical foundation of [12], where nonlinear maps are a priori assumed to be in the form of conditional certainty equivalents.
This result is obtained by solving an open question posed in [24] regarding the continuity (and joint measurability) of the state-dependent utility appearing in the representing functional of the preference relation (see the statement of Theorem 6 in Section 3). From a mathematical viewpoint Theorem 7 is the most triggering result, as it requires a quite structured and involved proof. Nevertheless such a result has the immediate benefit of clarifying the applicability of state-dependent preferences in optimization problems such as the maximization over a set of strategies.
1.1 Representation of time consistent nonlinear expectations
The representation of the conditional Chisini mean depicted so far allows to embed our theoretical results in the framework of nonlinear expectations. This deepens the understanding of the role of the Sure-Thing Principle in time consistent conditioning.
Kupper and Schachermayer [16] proved the noteworthy property that, under suitable structural assumptions, every time consistent family of strictly monotone, continuous and law invariant nonlinear maps take the form of a conditional certainty equivalent
$u^{-1}\left(\mathbb{E}_{\mathbb{P}}\left[u(X)\mid{\mathcal{F}}_{t}\right]\right),$ (1)
where $u$ is a strictly increasing function on a real interval. This result can be seen as a dynamic version of the aforementioned Nagumo-de Finetti-Kolmogorov Theorem.
A similar perspective can be observed in the recent stream of literature regarding the phenomenon of “collapse to the mean” for certain classes of functionals as in [3] for the law invariant case. Indeed the findings of our paper relate significantly to Theorem 6.1 in [10], where the assumption of law invariance is replaced with commonotonicity. In the realm of nonlinear expectations which are possibly not law invariant, the most important contribution comes from the theory of $g$-expectations related to Backward Stochastic Differential Equations, initiated and developed in [19, 8, 20]. The main drawback of such theory is that the notion of “conditionability” is constrained to the use of the Brownian filtration. Therefore in general it is not possible to find a conditional nonlinear expectation for finitely generated $\sigma$-algebras, that is also consistent with the considered $g$-expectation.
We overcome this issue by taking into consideration a non atomic [^1] Actually, requiring the existence of three disjoint sets in ${\mathcal{F}}$ with positive probability is sufficient to prove Theorem 2 probability space $(\Omega,{\mathcal{F}},\mathbb{P})$, the collection of sub-$\sigma$-algebras of ${\mathcal{F}}$
$\Sigma:=\left\{{\mathcal{G}}\middle|{\mathcal{G}}\text{ is a $\sigma$-algebra},\;{\mathcal{G}}\subseteq{\mathcal{F}}\right\},$
and a family of conditional nonlinear maps $\{{\mathcal{E}}_{{\mathcal{G}}}\}_{{\mathcal{G}}\in\Sigma}$, where ${\mathcal{E}}_{{\mathcal{G}}}:L^{\infty}(\Omega,\mathcal{F},\mathbb{P})\to L^{\infty}(\Omega,\mathcal{G},\mathbb{P})$, with $L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$ (resp. $L^{\infty}(\Omega,\mathcal{G},\mathbb{P})$) the quotient space of bounded random variables with respect to $\mathbb{P}$-almost sure equality. We denote with ${\mathcal{E}}_{0}$ the map obtained choosing ${\mathcal{G}}=\sigma(\emptyset)$, the trivial $\sigma$-algebra. Moreover we indicate by capital letters the elements in $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$, and by $\mathbf{1}_{A}$ the indicator functions [^2] Here we intend the equivalence class of such indicators. Later on we shall work also without quotienting adopting the same symbol to avoid heavy notations..
Definition 1.
We say that ${\mathcal{E}}_{{\mathcal{G}}}:L^{\infty}(\Omega,\mathcal{F},\mathbb{P})\to L^{\infty}(\Omega,\mathcal{G},\mathbb{P})$ is a conditional nonlinear expectation if for any $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$, $A\in{\mathcal{G}}$ we have ${\mathcal{E}}_{{\mathcal{G}}}(X\mathbf{1}_{A})={\mathcal{E}}_{{\mathcal{G}}}(X)\mathbf{1}_{A}$.
A family of conditional nonlinear expectations $\{{\mathcal{E}}_{{\mathcal{G}}}\}_{{\mathcal{G}}\in\Sigma}$ is time consistent if
${\mathcal{E}}_{0}({\mathcal{E}}_{{\mathcal{G}}}(X))={\mathcal{E}}_{0}(X)\text{for any }{\mathcal{G}}\in\Sigma\text{ and }X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P}).$ (2)
Assumption 1.
We assume that ${\mathcal{E}}_{0}$ is strictly monotone (on dicotomic random variables), i.e. for any $x,y\in\mathbb{R}$, $Z\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$ and $A\in{\mathcal{F}}$ with $\mathbb{P}(A)>0$
$x<y\text{ implies }{\mathcal{E}}_{0}(x\mathbf{1}_{A}+Z\mathbf{1}_{A^{c}})<{\mathcal{E}}_{0}(y\mathbf{1}_{A}+Z\mathbf{1}_{A^{c}}),$ (3)
and pointwise continuous, i.e. for any bounded sequence $X_{n}$ converging $\mathbb{P}$-almost surely to $X$
$\lim_{n\to\infty}{\mathcal{E}}_{0}(X_{n})={\mathcal{E}}_{0}(X).$ (4)
The following representation result can be seen as an enhanced version of (1) dropping the law invariance. Similarly to [16] and differently from Lemma 5.2 in [6], one of the most interesting aspects of our result is that the Sure-Thing Principle needs not to be explicitly assumed, but is rather deduced from the time consistency of the maps involved.
Theorem 2.
$\{{\mathcal{E}}_{{\mathcal{G}}}\}_{{\mathcal{G}}\in\Sigma}$ is a family of time consistent conditional nonlinear expectations such that ${\mathcal{E}}_{0}$ satisfies Assumption 1 if and only if  [^3] See the beginning of Section 5.5 for precise definitions
${\mathcal{E}}_{{\mathcal{G}}}(X)=\mathbb{u}_{\mathcal{G}}^{-1}(\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(X)\middle|{\mathcal{G}}\right])\quad\text{ for every }X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P}),$
for some $\mathbb{u}:\Omega\times\mathbb{R}\to\mathbb{R}$ which is jointly measurable, with $\mathbb{u}(\omega,0)=0$ and $\mathbb{u}(\omega,\cdot)$ continuous and strictly increasing for every $\omega\in\Omega$, $\mathbb{u}(\cdot,x)$ integrable for any $x\in\mathbb{R}$ and $\mathbb{u}_{\mathcal{G}}(\cdot,x)=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(x)\middle|{\mathcal{G}}\right]$.
1.2 Notations and preliminaries
@@
For any $f\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ and a sub $\sigma$-algebra ${\mathcal{G}}$, $\mathbb{E}_{\mathbb{P}}\left[f\middle|{\mathcal{G}}\right]$ is the conditional expectation i.e.
@@
In the following we shall assume familiarity with the properties of conditional expectation.
A preference relation is a binary relation $\succeq$ on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ which satisfies completeness and transitivity and induces a (strict) preference order $\succ$ by $f\succ g$ if and only if $f\succeq g$ but $g\nsucceq f$. In addition we set $g\sim f$ whenever $g\succeq f$ and $f\succeq g$. We denote by
@@
$\succeq$ satisfies the Sure-Thing Principle whenever, considering arbitrary $f,g,h\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and $A\in{\mathcal{F}}$ such that $f\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}\succeq g\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}$, we have $f\mathbf{1}_{A}+\tilde{h}\mathbf{1}_{A^{c}}\succeq g\mathbf{1}_{A}+\tilde{h}\mathbf{1}_{A^{c}}$ for every $\tilde{h}\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$;
@@
We say that $\succeq$ admits a numerical representation if there exists a functional $T:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ such that for any given $f,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$
$T(f)\geq T(g)\quad\text{ if and only if }\quad f\succeq g.$ (7)
2 Conditionable functionals and preferences
In [11] the notion of conditional Chisini mean was introduced as the solution of an infinite dimensional system of equations as we now recall in the following definition. Such a notion is intrinsically related to the structural properties of the functional $T$ involved, which in this paper will be the representing functional of $\succeq$.
Definition 2.
For a given functional $T:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$, we say that $T$ is conditionable on a sub $\sigma$-algebra ${\mathcal{G}}\subseteq{\mathcal{F}}$ if for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ there exists $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ such that
$T(f\mathbf{1}_{A})=T(g\mathbf{1}_{A})\quad\forall\,A\in{\mathcal{G}}.$ (8)
We say that $T$ is conditionable if it is conditionable on any ${\mathcal{G}}\in\Sigma$ and we call conditional Chisini mean $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ the family of solutions $g$ of Eq. (8).
If compared with (5), it evidently appears that $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ is the nonlinear counterpart of the standard conditional expectation. Although linearity is lost, nonlinear expectation still enjoys few useful properties of its linear counterpart such as, for instance, the “tower” and “taking out what is known” properties. Before the abstract formulation in [11], the problem of existence of nonlinear conditional expectation was already addressed in [8], yet the provided solution relies on a Brownian framework which in turn demands for more technical assumptions.
Embedding such a problem in a decision theoretical setup, we could rephrase the issue as follows: given a preference order $\succeq$ on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ then the certainty equivalent of $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ is $a\in\mathbb{R}$ such that $a\sim f$. Similarly if the agent is provided with additional information, represented by the $\sigma$-algebra ${\mathcal{G}}\subset{\mathcal{F}}$, then the ${\mathcal{G}}$-conditional certainty equivalent of $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ is any $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ such that
$f\mathbf{1}_{A}\sim g\mathbf{1}_{A}\quad\forall\;A\in{\mathcal{G}}.$ (9)
Loosely speaking, $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ can be considered as a ${\mathcal{G}}$-measurable counterpart of the ${\mathcal{F}}$-measurable function $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ (see also [12] and [17] for applications of the conditional certainty equivalent).
Since this paper deals with state-dependent preferences we shall directly refer to the conditional Chisini mean as the class
keeping in mind that Eq. (10) translates automatically to the framework in [11] as soon as we characterize a preference $\succeq$ through a representing functional (as defined in (7)).
Remark 3.
(SM) and (PC) of the preference $\succeq$ are sufficient to show the existence of a preference representing functional $T$ (see the proof of Proposition 9 for the further details).
We are now ready to state the first main result of this work, which links the Sure-Thing Principle to the conditionability of preferences and their representing functionals. The proof is postponed to Section 5.1.
Theorem 4.
Let $\succeq$ be a preference order satisfying (SM) and (PC). Then the following are equivalent
  1. (i)
    $\succeq$ satisfies (ST);
  2. (ii)
    any representing functional $T$ for $\succeq$ is conditionable;
  3. (iii)
    any representing functional $T$ for $\succeq$ is conditionable on [^4] Where $\sigma(A)=\left\{\Omega,\varnothing,A,A^{C}\right\}$ $\sigma(A)$ for any $A\in{\mathcal{F}}$.
We propose the following definition of regularity that is well suited for our purposes. In particular, it encompasses both the notion of continuity and that of joint-measurability that we seek to obtain for the state-dependent utility.
Definition 3.
For any ${\mathcal{G}}\in\Sigma$ we say that $\varphi:\Omega\times\mathbb{R}\to\mathbb{R}$ is a ${\mathcal{G}}$-regular function if
  • $\varphi(\omega,\cdot)$ is continuous and strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$;
  • $\varphi(\omega,0)=0$ for every $\omega\in\Omega$;
  • $\varphi$ is ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable.
Definition 4.
We say that representability holds for conditional Chisini means if there exist an ${\mathcal{F}}$-regular function $\mathbb{u}$, a probability measure $\mathbb{P}$ on ${\mathcal{F}}$ and, for every $\sigma$-algebra ${\mathcal{G}}\subseteq{\mathcal{F}}$, a ${\mathcal{G}}$-regular function $\mathbb{u}_{\mathcal{G}}$ such that $\mathbb{u}(\cdot,x)$ and $\mathbb{u}_{\mathcal{G}}(\cdot,x)$ are integrable for every $x\in\mathbb{R}$ and for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and $g\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ we can write [^5] More details can be found in the statement of Theorem 8.
$g(\omega)=\mathbb{u}_{\mathcal{G}}^{-1}(\omega,h(\omega))\quad\forall\,\omega\in\Omega,$
up to a choice of a version $h$ of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f)\middle|{\mathcal{G}}\right]$
Corollary 5.
Let $\succeq$ be a preference order satisfying (SM) and (PC). Whenever ${\mathcal{F}}$ contains at least three disjoint non-null events, condition (i) of Theorem 4 is equivalent to
  1. (iv)
    conditional Chisini means are representable.
3 Solution to an open question in Wakker and Zank (1999)
The proofs of the results stated in the previous sections mostly rely on an enhanced version of the preference representation result due to Wakker and Zank [24], which we are now going to present in detail. For a general state space $(\Omega,{\mathcal{F}})$, Wakker and Zank [24] provide a characterization of those preference relations $\succeq$ on the space of bounded random acts which admit a numerical representation through an additively decomposable functional ([24], Theorem 11). This result extends to an infinite dimensional setup the theory developed by Debreu [9]. By adopting the stronger notion of pointwise continuity of $\succeq$, Theorem 11 in [24] can be specialized to the following integral representation form.
Theorem 6 ([24], Theorem 12 and [5], Theorem 5).